OpenStudy (anonymous):

Find the indicated limit, if it exists. limit of f of x as x approaches 0 where f of x equals 5 x minus 8 when x is less than 0 and the absolute value of the quantity negative 4 minus x when x is greater than or equal to 0

OpenStudy (dan815):

f(x)=5x-8 } x<0 f(x)=4-x } x>0

OpenStudy (dan815):

like that?

OpenStudy (dan815):

oh -4 sry

OpenStudy (dan815):

f(x)=5x-8 } x<0 f(x)=|-4-x| } x>0

OpenStudy (dan815):

the limt DNE

OpenStudy (dan815):

since the value f(0-) != f(0+)

OpenStudy (dan815):

the values left of zero are not equal to the value at just right of zero

OpenStudy (anonymous):

okay these are y answer choices -4 -12 -8 The limit does not exist. @dan815

OpenStudy (dan815):

okay so these are the rules

OpenStudy (dan815):

the limit doesnt exist at the point x=0 if from the left size of 0 you approach a different number as from left side

OpenStudy (solomonzelman):

you mean that if from one side of 0 the limit approaches one number and from the other side of 0 the limit approaches a different number, than the limit doesn't exist.